OpenStudy (anonymous):

4/m-5 - 3/m+5 = 2/m^2-25

OpenStudy (anonymous):

denominator: (m-5)(m+5) 4/(m-5) --> make denominator (m-5)(m+5) ---> 4(m+5)/(m-5)(m+5) 3/(m+5) --> make denominator (m-5)(m+5)---> 3(m-5)/(m-5)(m+5) (4m+20-3m-15)/(m-5)(m+5) = m+5/(m-5)(m+5) = 1/(m-5)

OpenStudy (radar):

m=-33

OpenStudy (anonymous):

ohh sorry i got the question wrong

OpenStudy (radar):

No problem, I think I got the answer wrong!

OpenStudy (radar):

suzi20 i have been looking at your solution, and I believe the last line should be: (4m+20+3m-15)/(m-5)(m+5) or (7m+5)/(m^2-25) since this is equal to 2/(m^2-25) we can multiply both sides by (m^2-25) and get: 7m+5=2 7m=-3 m=-3/7

OpenStudy (radar):

I think that is better than the -33.

OpenStudy (anonymous):

thats not one of the answers to choose from radar

OpenStudy (anonymous):

im exhausted i give up

OpenStudy (anonymous):

the question is to proof 4/m-5 - 3/m+5 = 2/m^2-25 or to get the value of m?

OpenStudy (anonymous):

yes suzi

OpenStudy (anonymous):

proof or value?

OpenStudy (radar):

I give up too

OpenStudy (radar):

was the answer m=-33?

OpenStudy (radar):

Come on top dollar answer the questions

OpenStudy (radar):

Were you supposed to solve for m?

OpenStudy (anonymous):

radar im not sure but thats what im putting

OpenStudy (radar):

Sorry I couldn't of been more help. It's late for me maybe tomorrow. I am writing the problem down.

OpenStudy (anonymous):

ok thanks and good night i gotta study for my history exam

OpenStudy (radar):

Good morning topdollar843. I'm gonna try again (lol) 4/(m-5) -3/(m+5) = 2/(m^2-25). Common denominator is (m^2-25) 4(m+5/(m^2-25) - 3(m-5)/(m^2-25)= 2/(m^2-25) 4m+20-3m+15/(m^2-25) = 2/(m^2-25) consolidating similar terms (m+35)/(m^2-25)=2/(m^2-25) multiplying both sides by (m^2-25) m+35=2 m=-33

OpenStudy (radar):

I got this the first time but I was so sloppy in my work, I didn't trust it! I feel pretty good that m is -33